OpenStudy (anonymous):

Write the product in standard form. ( 7 + 7i)( 6 + 7i)

OpenStudy (anonymous):

( 7 + 7i)( 6 + 7i) 42+49i+42i+49i^2 49i^2+91i+42

OpenStudy (anonymous):

remember than i^2 = -1 . must be simplified

OpenStudy (anonymous):

Use distributive law (a+b) (c+d) =ac+ad+bc+bd) here a=7, b=7i=d ,c=6 So the solution will be 7*6+7*7i+6*7i+7i*7i =42+49i+42i+49*i^2 (Since i^2= -1) we have =42+91i-49 =(-7+91i) will be the Answer

OpenStudy (anonymous):

Really @Anu2401 ??? :o

OpenStudy (anonymous):

@Anu2401 Thank you!